Imagine you are 15 years old. Your father is 45. In how many years will your father be twice as old as you? ( algebraic expression needed please and thank you)
Answer by mananth(16946)

My age 15 years
Fathers age 45 years
difference = 30 years
let my father be twice my age after n years
my age after n years 15+n
Father's age after n years
45 +n
SO
45+n = 2(15+n)
45+n = 2n +30
n =15
after 15 years
15+15 = 30
15+45 = 60